What is the angle at which both the height and the range of projectile is maximum?
1
Expert's answer
2019-12-25T12:13:53-0500
"R=\frac{v^2}{g}\sin{2\theta}=\frac{v^2}{g}2\sin{\theta}\cos{\theta}"

"H=\frac{v^2}{2g}\sin^2{\theta}"

So,

"\frac{v^2}{g}2\sin{\theta}\cos{\theta}=\frac{v^2}{2g}\sin^2{\theta}"

"\tan{\theta}=4"

"\theta=\arctan{4}=76\degree"